### v1

# Best of both worlds: Fusing hyperspectral data from two generations of spectro-imagers for X-ray astrophysics

Abstract
Context. With the recent launch of the X-Ray Imaging and Spectroscopy Mission (XRISM) and the advent of microcalorimeter detectors, X-ray astrophysics is entering in a new era of spatially resolved high resolution spectroscopy. But while this new generation of X-ray telescopes have much finer spectral resolutions than their predecessors (e.g. XMM-Newton, Chandra), they also have coarser spatial resolutions, leading to problematic cross-pixel contamination. This issue is currently a critical limitation for the study of extended sources such as galaxy clusters of supernova remnants.
Aims. To increase the scientific output of XRISM’s hyperspectral data, we propose to fuse it with XMM-Newton data, and seek to obtain a cube with the best spatial and spectral resolution of both generations. This is the aim of hyperspectral fusion. In this article, we explore the potential and limitations of hyperspectral fusion for X-ray astrophysics.
Methods. We have implemented an algorithm that jointly deconvolves the spatial response of XRISM and the spectral response of XMM-Newton. To do so, we construct a forward model adapted for instrumental systematic degradations and Poisson noise, then tackle hyperspectral fusion as a regularized inverse problem. We test three methods of regularization: spectral low rank approximation with a spatial Sobolev regularization; spectral low rank approximation with a 2D wavelt sparsity constraint ; and a 2D-1D wavelet sparsity constraint.
Results. We test our method on toy models constructed from hydrodynamic simulations of supernova remnants. We find that our method reconstructs the ground truth well even when the toy model is complex. For the regularization term, we find that while the low rank approximation worked well as a spectral denoiser in models with less spectral variability, it introduced a bias in models with more spectral variability, in which case the 2D-1D wavelet sparsity regularization worked best. After demonstrating a proof of concept in this article, we aim to apply this method to real X-ray astrophysical data in the near future.
Key Words.:
Methods: data analysis, X-rays: general
1 Introduction
With the advent of microcalorimeter detectors, X-ray astrophysics has entered in a new era of spatially resolved high resolution spectroscopy. Indeed, the recent launch of the X-ray observatory XRISM and its microcalorimeter array Resolve [^1] https://xrism.isas.jaxa.jp, and the future Athena telescope and its XIFU [^2] https://x-ifu.irap.omp.eu/ camera, are beginning to grant access to spatially resolved spectroscopy at spectral resolutions improved by an order of magnitude. This will allow astrophysicists to investigate previously indistinguishable spectral features, such as the emission lines of rare elements, line broadening, or velocity measurements. These characteristics are essential in the study of extended sources such as supernova remnants, galaxy clusters, and many of the other exotic sources of the X-ray universe.
However, while the XRISM/Resolve instrument has a much finer spectral resolution than Chandra/ACIS [^3] https://cxc.harvard.edu/ or XMM-Newton/EPIC [^4] https://www.cosmos.esa.int/web/xmm-newton cameras, its spatial resolution is much lesser than that of its predecessors. As we will see, this spatial blur is an important obstacle when analyzing XRISM/Resolve data. Given this trade-off, could one combine the data from these two generations of X-ray observatories and obtain a result with the best spectral and spatial resolutions of each instrument? This is the idea at the heart of hyperspectral fusion. In this paper, we explore the potential and limitations of hyperspectral fusion for X-ray astrophysics, and test our method on simulations resembling a supernova remnant’s observations.
Figure 1: Comparison of hyperspectral data of the Perseus galaxy cluster as observed by two generations of X-ray spectro-imagers. The XMM-Newton/MOS (top) and Hitomi/SXS (bottom) data have pixels of 8”/30” and spectral channels of 15/1 eV respectively. The left panels show the counts maps summed over all spectral channels and the right panels the spectra summed over all pixels.
1.1 Astrophysical context: X-ray spectro-imaging and microcalorimeters
X-ray spectro-imagers can collect hyperspectral images. These are three-dimensional data arrays: they have two spatial dimensions, like a classical image, and a third spectral dimension (energy, or wavelength). That is to say, to each spatial pixel is associated an energy spectrum.
In the last decades, Charge-coupled device (CCD) cameras have been the main spectro-imaging instruments onboard X-ray telescopes such as XMM-Newton, Chandra, Swift, Suzaku and eROSITA. When the X-ray photon interacts with the semiconductor material, an electric charge is generated proportional to the energy of the incident photon. The charge is then transferred row by row to a readout register. For example, the XMM-Newton MOS camera can accommodate a large number (600$\times$600) of small pixels (40 $\mu$m, Turner et al., 2001). The energy resolution is of the order 150 eV [^5] https://xmm-tools.cosmos.esa.int/external/xmm_user_support/documentation/uhb/epic_specres.html at 6 keV ($E\,/\Delta E\sim 40$) which is close to the Fano limit, the physical lower bound of energy resolution due to the statistical fluctuations in the number of electron-hole pairs produced in the photon-material interaction.
To overcome this limit, a change of technology is required. Microcalorimeter devices are providing this revolutionary leap forward for X-ray spectro-imaging instruments. The fundamental principle of microcalorimeters is to measure the photon energy via the small increase in temperature caused by the conversion of this energy into heat in the sensor. To be sensitive to small heat increments and reach a spectral resolution in the order of a few eV, the device has to be cooled down to 50 mK with very stable temperature. It is important to note that, unlike gratings onboard XMM-Newton or Chandra, microcalorimeters are non-dispersive spectrometers and can be used to study diffuse structure.
However, with the current available technology, only a limited number of pixels can be accommodated on the focal plane. The XRISM/Resolve instrument is a microcalorimeter array of 6 × 6 sensors (pixel size of 800 $\mu$m, Takahashi et al., 2014) with a constant spectral resolution over the entire bandpass of 5 eV FWHM (XRISM Science Team, 2020). While future X-ray telescopes combining microcalorimeter and high spatial resolution are planned (Athena/XIFU, see Barret et al., 2018, for more details), they are expected to launch in the late 2030’s, and would still not reach the spatial resolution of the X-ray telescope Chandra.
In the meantime, mitigation plans to overcome this spatial resolution limitation should be investigated to try to partially recover some of the spatial information. To that end, we explore in this study the performance of hyperspectral fusion to combine data sets with high spectral resolution and high spatial resolution, using the example of fusion between XMM-Newton and XRISM/Resolve data. The final goal of such reconstruction is to enhance the scientific return of XRISM/Resolve by retrieving maps of physical parameters (plasma temperature, velocity, abundances, etc) at higher resolution with reduced cross-contamination betweeen pixels (spectrospatial mixing).
Indeed, this mixing is an important problem when analysing hyperspectral data. If an instrument’s spatial resolution is larger than the coherence scale of a physical parameter, the reconstructed information will be an average of the small scale features. To take the example of a velocity map from the ejecta of a supernova remnant, the small scale redshifted/blueshifted ejecta features can mistakenly be measured at small or close to zero redshift values, as was shown in Godinaud et al. (2023)’s detailed analysis of Tycho’s supernova remnant. By jointly deconvolving the data sets from two X-ray instruments, hyperspectral data fusion could solve this problem and recover a higher fidelity scientific information at small scale.
1.2 Hyperspectral Fusion
Given two hyperspectral (HS) images, the general task of hyperspectral fusion consists in obtaining a cube with the best spatial and spectral resolution of each data set. For instance, Fig. 1 shows HS images of the Perseus galaxy cluster taken respectively by the X-ray telescopes XMM-Newton and Hitomi/SXS [^6] XRISM/Resolve camera is identical to the SXS onboard the short-lived Hitomi telescope which was launched in 2016.. XMM-Newton has better spatial resolution, and Hitomi has better spectral resolution. The goal of fusing the two datasets would be to obtain a HS image with XMM-Newton’s spatial resolution and Hitomi’s spectral resolution.
This problem has been explored in the literature for applications to satellite pictures of the earth, and has recently received some attention in astrophysics with the launch of the James Webb Space Telescope (JWST). However, X-ray astrophysics has its own specific challenges. X-ray data is dominated by Poisson noise, and both instruments produce HS images with many spectral channels, as opposed to one of the data sets being a multispectral (MS) image with a small amount of spectral bands typically obtained with different filters. To use hyperspectral fusion for X-ray astrophysics, it is thus key to find the fusion method most appropriate for this case study.
In this section, we give an overview of the various aspects to take into account when approaching a problem of hyperspectral fusion, in order to find the best way to solve the problem. For a detailed review of the literature, we refer the interested reader to Yokoya et al. (2017).
The first aspect to consider is the forward model of the two instruments. This defines how a source image will be degraded in the process of image formation and recording. It will depend on the systematic characteristics of the instrument, and the noise we expect to measure.
For the systematic instrumental degradation, two dimensions may be taken into account: the spatial degradation, and the spectral degradation. Spatially, the degradation will be due to optical diffraction or other systematic blurs, and this information is encapsulated in the point source function (PSF). Given a point source, the PSF returns its radiance distribution on the recorded image. Similarly, if a source emits a Dirac spectra, the measured spectra will be the instrument’s spectral response function, and this will be affected by, for instance, the chemical and physical properties of the spectro-imager. Depending on the instrument, the PSF and spectral response may depend on the energy, on the spatial location of the pixel on the camera, or both.
Then, for the stochastic noise, while most of the literature on fusion assumes additive Gaussian noise, it is important to realize this is not always appropriate. In X-ray astrophysics, photons have a low flux and are measured individually, and thus follow a Poisson distribution.
Once the instrumental systematic degradation and the stochastic noise have been taken into account, if we take two HS images $X$ and $Y$ of sizes ($M\times N\times l$) and ($m\times n\times L$), the general problem to fuse two HS images will be:
$\hat{Z}=\operatorname*{arg\,min}_{Z}D(\mathcal{M}_{X}(Z)|X)+D(\mathcal{M}_{Y}(Z)|Y),$ (1)
where $\hat{Z}$ is the fusion result, $\mathcal{M}_{X}$ and $\mathcal{M}_{Y}$ are the model operators that take into account the systematic instrumental degradation (PSF, spectral response…) and the rebinning needed to take account different voxel sizes, and $D$ is a well-chosen function that evaluates the difference between the model $\mathcal{M}_{X}(Z)$ and data $X$, and between model $\mathcal{M}_{Y}(Z)$ and data $Y$, given the expected noise distribution. Thus, this problem ensures that $Z$ is a plausible source from which the first instrument would measure $X$, and the second would measure $Y$. In the case of X-ray astrophysics, the forward model is the one described in section 2.1 and Fig. 2.
However, this problem is often ill-posed (in the sense of Hadamard, 1923) and ill-conditioned, meaning that while there might be a unique solution, the noise will be amplified when inverting the model, corrupting the solution, and slight changes in the data would lead to drastic changes in the solution (Bertero M. & C., 2022). One well-studied approach to solve ill-posed inverse problems is to regularize the problem, which restricts the space of acceptable solutions. This can be approached by considering the spatial features of $Z$, its spectral features, or both. To regularize in the spectral dimension, one classic method is to project the spectra of $Z$ onto a lower-dimensional subspace, such as is done with source separation (e.g. Yokoya et al., 2012; Prévost et al., 2022; Pineau et al., 2023) or low-rank approximation (e.g. Simões et al., 2015; Xu et al., 2022; Guilloteau et al., 2020). We can also add a regularization term $\varphi(Z)$ to the function we seek to minimize, which leads us to:
$\hat{Z}=\operatorname*{arg\,min}_{Z}D(\mathcal{M}_{X}(Z)|X)+D(\mathcal{M}_{Y}(Z)|Y)+\varphi(Z).$ (2)
This $\varphi(Z)$ will promote certain spatial or spectral structures in the solution. Its choice must thus be appropriate for our data. For instance, in remote sensing (i.e. the study of satellite pictures of the earth), it may be useful to promote spatial piece-wise smoothness using a total variation term such as done by Simões et al. (2015), but this would be ill-advised for astrophysical data, where such features are rarely observed. For our case of X-ray astrophysics, we consider three potential regularization approach, which are described in section 2.3.
Finally, to solve the problem in equation 2, we need to pick an appropriate optimization scheme. In some cases, there might be an analytical solution (such as Pineau et al. (2023)’s work for JWST, when the spectral content is assumed to be a linear combination of known spectra), but most often an iterative algorithm will have to be constructed. The choice of optimizer will depend on the data size, the specific challenges presented by the forward model, and the chosen regularization terms. It may also be the case that constraints placed on the optimizer (such as, wishing that the algorithm reaches convergence as quickly as possible) will lead to constraints on the choice of regularization and cost function.
Figure 2: A simplified diagram of the forward model of two X-ray telescopes (here, XMM-Newton/EPIC and XRISM/Resolve). The problem of fusion consists in inverting this model, using the observed data X and Y to retrieve the deconvoluted cube Z. The mathematical details behind the operators and kernels are detailed in section 2. Not shown is the exposure time of each data sets.
1.3 Contributions
In astrophysics, work has been done to tackle MS/HS fusion for the JWST (see Guilloteau et al., 2022; Pineau et al., 2023), to fuse images taken by two of its instruments (NIRCam and NIRSpec), but to our knowledge, no algorithm has been implemented for HS/HS fusion with Poisson statistics, considering two instruments with their PSF and spectral responses, as is needed in X-ray astrophysics.
In this work, we investigate a hyperspectral fusion method for X-ray astrophysics. Our aim is to evaluate the potential and limitations of fusion for X-ray data taken by couples of telescopes such as XMM-Newton/EPIC with XRISM/Resolve, or Chandra with Athena-XIFU. Section 2 presents our methodology, going over the forward model we have developed for X-ray astrophysics, and the three potential regularization schemes we consider: first, inspired by Guilloteau et al. (2020), and given the widespread usage of low-rank approximations in the literature, we implement a low-rank approximation with Sobolev spatial regularization. Since Sobolev regularization does not take into account multi-scale features, we then consider a low-rank approximation with a 2D wavelet sparsity regularization. Finally, we investigate whether low-rankedness is indeed appropriate, and also implement a heuristic approach using a 2D-1D Wavelet sparsity regularization.
We have implemented a modular optimization algorithm which can use either of these three potential regularizations. Its code is openly available on Github. [^7] https://github.com/JMLascar/HIFReD_fusion In section 3, we test our method on toy models of varying complexity, using synthetic X-ray data cubes constructed from hydrodynamic simulation of supernova remnants. We discuss the achievements, limitations, and perspective for future work in section 4.
2 Methodology
In this section, we present the methodology of our fusion algorithm for X-ray astrophysics. We go over the general problem at play in section 2.1. Section 2.2 goes over the rebinning operator needed when the data sets have different voxel sizes. Section 2.3 covers our choices of potential regularization terms. Section 2.4 goes over our algorithmic strategy to solve the inverse problem.
2.1 Forward Model
Let $X$ be the hyperspectral data from the instrument with better spatial resolution, but worse spectral resolution (e.g. XMM-Newton/EPIC). Let $Y$ be the data from the other instrument, with worse spatial resolution and better spectral resolution (e.g. XRISM/Resolve). When available, $\bar{{X}}$ and $\bar{{Y}}$ are the ground truths at the respective instruments’ resolutions.
In this section, we consider the problem when both HS images have the same size, i.e. $X$ and $Y$ +$\in\mathbb{R}^{M\times N\times L}_{+}$, and thus so does $Z$ (for the case where $X$ and $Y$ have different sizes, see section 2.2). Our forward model (seen in Fig. 2 ) must include each instrument’s spatial degradation, which is modelled by their PSFs, as well as the spectral degradation, modelled by their spectral responses. We must also account for the effective area of the instrument, and the exposure time. For the former, we assume an effective area with no vignetting, i.e. the effective area does not vary spatially. [^8] Accounting for vignetting would be easy, however, since the effective area is simply multiplied to the model term by term. Thus whether we consider the effective area as a spatially uniform cube or a vignetted cube makes no difference to the algorithmic complexity. We call $A_{X}$ and $A_{Y}$ the effective area cubes, which have the same shape as $X$ and $Y$. $t_{X}$ and $t_{Y}$ are scalars, the exposure times of each data set. The instruments’ spectral responses are $\textbf{s}_{X}$ and $\textbf{s}_{Y}$, and their PSFs are $B_{X}$ and $B_{Y}$.
We define $\mathcal{B}_{X}$ and $\mathcal{B}_{Y}$ as the operators which, given a cube $Z$, apply a 2D convolution on its spatial slices with the PSF $B_{X}$ or $B_{Y}$. Similarly, we define as $\mathcal{S}_{X}$ and $\mathcal{S}_{Y}$ the operators which convolve each spectra of $Z$ with $\textbf{s}_{X}$ or $\textbf{s}_{Y}$. Formally, if $Z_{k}$ is the 2D slice of $Z$ at the $k$th spectral channel, and $Z_{ij}$ is the spectra of $Z$ at pixel $(i,j)$:
$\Big{(}\mathcal{B}_{Y}Z\Big{)}_{k}=B_{Y}\circledast Z_{k}$ (3)
$\Big{(}Z\mathcal{S}_{X}\Big{)}_{ij}=Z_{ij}\circledast S_{X},$ (4)
[…]
The aim is to reach, as best as possible, a fused cube $Z$ that has the spatial resolution of $X$ and the spectral resolution of $Y$. The general problem is:
$\hat{Z}=\operatorname*{arg\,min}_{Z\geq 0}\mathcal{L}_{P}\Big{(}X\;|\;Z_{X}\Big{)}+\mathcal{L}_{P}\Big{(}Y\;|\;Z_{Y}\Big{)}+\varphi(Z),$ (5)
where:
$Z_{X}=(t_{X}Z\odot A_{X})\mathcal{S}_{X}$
$Z_{Y}=\mathcal{B}_{Y}(t_{Y}Z\odot A_{Y}),$ (6)
and the data-fidelity metric $\mathcal{L}_{P}$ is the negative Poisson log-likelihood, $\varphi(Z)$ is the regularization term. We impose $Z\geq 0$ as there is no physical interpretation for a negative spectra. The likelihood terms promote faithfulness between our model and the data at comparable resolutions.
It is worth noting that in order to fully invert the forward model, we would need to compare $X$ with $\mathcal{B}_{X}(t_{X}Z\odot A_{X})\mathcal{S}_{X}$, and likewise for $Y$. However, we are not interested in retrieving a cube $Z$ with infinite resolution, but rather wish $Z$ to have the spatial resolution of $X$, and the spectral resolution $Y$. The problem expressed in Equations 5 and 6 will promote such a solution, since no spatial degradation is applied to go from $Z$ to $Z_{X}$, and likewise spectrally for $Z_{Y}$ with $Y$.
More explicitly, we obtain:
$\hat{Z}=$ $\operatorname*{arg\,min}_{Z\geq 0}\sum_{i=1}^{M}\sum_{j=1}^{N}\sum_{k=1}^{L}\Bigg{[}\Big{(}(t_{X}Z\odot A_{X})\mathcal{S}_{X}\Big{)}_{ijk}$
$-X_{ijk}ln\Big{(}(t_{X}Z\odot A_{X})\mathcal{S}_{X}\Big{)}_{ijk}$
$+\Big{(}\mathcal{B}_{Y}(t_{Y}Z\odot A_{Y})\Big{)}_{ijk}-Y_{ijk}ln\Big{(}\mathcal{B}_{Y}(t_{Y}Z\odot A_{Y})\Big{)}_{ijk}\Bigg{]}+\varphi(Z),$ (7)
where the shape of $Z$, $X$, $Y$, are all $(M,N,L)$.
2.2 Rebinning Operator
In most cases, two different hyperspectral instruments will not produce data of the same voxel size. Rather, we will have $X\in\mathbb{R}_{+}^{M\times N\times l}$, and $Y\in\mathbb{R}_{+}^{m\times n\times L}$, where $L>l$, $M>m$, and $N>n$. In other words, $X$has more pixels than $Y$, and $Y$ has more spectral channels than $X$. Our result $Z$ should be of size $(M\times N\times L)$. In order to compare $Z$ to the data $X$ and $Y$ in the data fidelity terms of equation 5, we thus need to change the size of $Z$ to that of $X$ and $Y$ respectively. This is done by applying a rebinning operator.
For the sake of simplicity, and to keep this operator differentiable, we assume that $M$ is a multiple of $m$, $N$ of $n$, and $L$ of $l$. Let us call the multiplying factors $(\nu_{M},\nu_{N},\nu_{L})$, such that $M=\nu_{M}m$, and likewise for the other dimensions. We need three rebinning operators: $\mathcal{R}_{M}$, $\mathcal{R}_{N}$, and $\mathcal{R}_{L}$.
In this case, the rebinning operator is that which sums together pixels in nearby groups of $\nu$.
For instance, for $\mathcal{R}_{M}$, a single voxel $(s,j,k)$ of the rebinned tensor is:
$\big{(}\mathcal{R}_{M}\textbf{Z}\big{)}(s,j,k)=\sum_{p=s*\nu_{M}}^{p=(s+1)*\nu_{M}-1}\textbf{Z}(p,j,k)$ (8)
Such an operator can be neatly written as a matrix product between each slice of the tensor $Z$ and a block diagonal matrix filled only with zeroes and ones. For simplicity, we’ll denote as $\mathcal{R}_{MN}$ the spatial rebinning operator (which applies $\mathcal{R}_{M}$ on the lines, and $\mathcal{R}_{N}$ on the columns).
Now, to include rebinning in our optimization problem posed in equation 5, $Z_{X}$ and $Z_{Y}$ become:
$Z_{X}=(t_{X}Z\odot A_{X})\mathcal{S}_{X}\mathcal{R}_{L}$
$Z_{Y}=\mathcal{R}_{MN}\mathcal{B}_{Y}(t_{Y}Z\odot A_{Y}),$ (9)
and plugging this in the Poisson likelihood:
$\hat{Z}=$ $\operatorname*{arg\,min}_{Z\geq 0}\sum_{i=1}^{M}\sum_{j=1}^{N}\sum_{k=1}^{L}\Bigg{[}\Big{(}(t_{X}Z\odot A_{X})\mathcal{S}_{X}\mathcal{R}_{L}\Big{)}_{ijk}$
$-X_{ijk}ln\Big{(}(t_{X}Z\odot A_{X})\mathcal{S}_{X}\mathcal{R}_{L}\Big{)}_{ijk}+\Big{(}\mathcal{R}_{MN}\mathcal{B}_{Y}(t_{Y}Z\odot A_{Y})\Big{)}_{ijk}$
$-Y_{ijk}ln\Big{(}\mathcal{R}_{MN}\mathcal{B}_{Y}(t_{Y}Z\odot A_{Y})\Big{)}_{ijk}\Bigg{]}+\varphi(Z).$ (10)
2.3 Regularization
In X-ray astrophysics, our main issue will be the low signal-to-noise ratio, as HS images are dominated by Poisson noise. Thus, we are looking for regularization terms that will be efficient denoisers for Poisson statistics, and which will be appropriate for astrophysical images (often displaying diffuse and/or isotropic features) and X-ray spectra (which can present strong spectral variability).
Our first consideration was sparsity under a 2D-1D starlet transform. Then, inspired by Guilloteau et al. (2022)’s work for JWST, we implemented a low rank approximation with Sobolev spatial regularization. Finally, we wished to test a compromise between the two, with a low rank approximation and a 2D spatial wavelet sparsity regularization. The following sections describe these three options.
A 2D-1D Wavelet Sparsity
A classic regularisation method is to choose a term that promotes the sparsity of $Z$ under a well-chosen orthogonal transform — that is to say, that in some domain, $Z$ can be represented by only a small amount of non-zero coefficients. Wavelets are useful to capture multi-scale features, and in particular, the starlet transform is appropriate for astrophysical data due to its isotropic structure (Starck & Murtagh, 1994).
More specifically, we chose the 2D-1D Starlet transform, which we define as applying the two-dimensional Starlet transform channel-wise, then the one-dimensional Starlet transform pixel-wise. Thus, in this case $\varphi(Z)$ is the $l_{1}$ norm (the minimization of which promotes sparsity, see Bach et al., 2011) of the 2D-1D Starlet transform coefficients of $Z$. For $S$ wavelet scales, we define the 2D-1D starlet transform as $\mathcal{W}_{S}^{2D1D}$, and the $l_{1}$ norm regularization term is:
$\varphi(Z)$ $=\Big{|}\Big{|}\mathcal{W}_{S}(Z)\Big{|}\Big{|}_{1}$
$=\mu\sum_{i=1}^{M}\sum_{j=1}^{N}\sum_{k=1}^{L}\sum_{s=1}^{S}|\mathcal{W}_{S}(Z)^{2D1D}_{i,j,k,s}|,$ (11)
where $\mu$ is a hyperparameter to control the intensity of regularization.
B Low rank and Sobolev
This regularization option was inspired by Guilloteau et al. (2020)’s work for JWST, though it was implemented from scratch so it could be adapted for HS-HS fusion and the specifics of X-ray astronomy. The spectral regularization is a low-rank approximation, which corresponds to decomposing $Z=WV,$ where $V\in\mathbb{R}^{L\times r}$ spans the subspace of rank $r$ to project $Z$ onto, and $W\in\mathbb{R}^{r\times M\times N}$ is the tensor of representation coefficients in that subspace. $V$ is found by performing principal component analysis on $Y$ (the hyperspectral image with the most spectral information) and keeping only the highest singular values. So now $Z_{X}$ and $Z_{Y}$ from Equation 9 become variables of $W$:
$Z_{X}(W)=(t_{X}WV\odot A_{X})\mathcal{S}_{X}\mathcal{R}_{L}$
$Z_{Y}(W)=\mathcal{R}_{MN}\mathcal{B}_{Y}(t_{Y}WV\odot A_{Y}),$ (12)
And can be plugged into Equation 10, now seeking to minimize $W$. For spatial regularization, the Sobolev regularization (Karl, 2005) implies adding the following term to the cost function:
$\varphi(W)=\mu||\mathcal{D}(W)||^{2}_{F},$ (13)
where $\mathcal{D}$ is a first order 2-D finite difference operator. This term will thus aim to minimize the differences between neighbouring pixels on the representation coefficients of the subspace spanned by $V$.
Both the Sobolev regularization and the low-rankedness will result in a faster algorithm, due to the dimension reduction and simple expression of the $\mathcal{D}$ operator. However, Sobolev does not capture multi-scale spatial features. Further, since $V$ is extracted from $Y$, it may reproduce features found in $Y$ too strongly. This is fine (and even desirable) in $MS/HS$ fusion, where the MS image does not bring much spectral information, but in $HS/HS$ fusion, $X$ contains spectral information, in particular because it displays much less spatial mixing than $Y$. The assumption that all of $Z$’s spectra can be decomposed as linear combinations of $Y$ is not necessarily true, and forcing a low-rank approximation in this case will bring on spectral distortions. If this assumption is sound, however, low-rankedness will be an undeniable asset.
C Low rank and 2D Wavelet Sparsity
Since Sobolev regularization does not capture multi-scale information the way wavelets do, we considered the option to keep a low-rank approximation as described in the previous section, then to use a starlet sparsity term for the spatial regularization:
$\varphi(W)=\mu||\mathcal{W}_{S}^{2D}(W)||_{1},$ (14)
where the operator $\mathcal{W}_{S}^{2D}$ applies the 2D starlet transform to every slice of $W$.
2.4 Algorithm Description
A Proximal Gradient Descent
If we consider the problem posed in Equations 7 or 10, and replace the regularization term by one of the three options we just considered, we notice that the cost function is convex, and can be separated in a differentiable and non-differentiable part. In such cases, proximal gradient methods have been shown to be very effective (Parikh & Boyd, 2014).
The proximal gradient descent is an iterative method where, for each iteration, we first undergo a gradient descent step on the differentiable part of the cost function, then we apply a proximal operator (prox), the purpose of which is to find a solution close to the result obtained by gradient descent while minimizing the function’s non-differentiable part. Formally, if the result obtained by the gradient descent step is $x$, and the non-differentiable part of the function is $\varphi$, the operator is:
$\textbf{prox}_{\varphi}(x)=\operatorname*{arg\,min}_{y}(\varphi(y)+\frac{1}{2}||y-x||_{2}^{2}).$ (15)
For the $l_{1}$ norm, the proximal operator is a soft thresholding function:
$\textbf{prox}_{l_{1},\mu}(x)=\cases{0}&\text{if }|x|<\mu\\ x-\mu\cdot sign(x)\text{if }|x|\geq\mu\\ {},$ (16)
where $\mu$ is the parameter that controls the strength of regularization. For the non-negativity constraint, the prox is simply the operator which puts all negative values to 0.
The gradient of the differentiable part in Equation 5 is:
$\nabla_{Z}\mathcal{L}_{P}$ $=\nabla_{Z}\Big{(}\mathcal{L}_{P}(X\;|\;Z_{X})\Big{)}+\nabla_{Z}\Big{(}\mathcal{L}_{P}(Y\;|\;Z_{Y})\Big{)}$
$=\frac{\partial(Z_{X})}{\partial{Z}}\nabla_{Z_{X}}\Big{(}\mathcal{L}_{P}(X\;|\;Z_{X})\Big{)}+\frac{\partial(Z_{Y})}{\partial{Z}}\nabla_{Z_{Y}}\Big{(}\mathcal{L}_{P}(Y\;|\;Z_{Y})\Big{)}$ (17)
In the case with rebinning, we have:
$\frac{\partial(Z_{X})}{\partial{Z}}=t_{X}A_{X}\mathcal{R}_{L}^{\dagger}\mathcal{S}_{X}^{\dagger},$
$\frac{\partial(Z_{Y})}{\partial{Z}}=t_{Y}A_{Y}\mathcal{B}_{Y}^{\dagger}\mathcal{R}_{MN}^{\dagger},$ (18)
where $\dagger$ denotes the conjugate transpose of the operators, in the sense of, if we were to vectorize the spatial dimensions of our data and write the operator $\mathcal{S}$ as a product with the 2D matrice $S$, $\mathcal{S}^{\dagger}$ would apply the product with matrix $S^{\dagger}$. In the case with a low rank approximation, where the gradient descent is done on $W=ZV^{T}$:
$\frac{\partial(Z_{X})}{\partial{W}}=Vt_{X}A_{X}\mathcal{R}_{L}^{\dagger}\mathcal{S}_{X}^{\dagger},$
$\frac{\partial(Z_{Y})}{\partial{W}}=Vt_{Y}A_{Y}\mathcal{B}_{Y}^{\dagger}\mathcal{R}_{MN}^{\dagger}.$ (19)
In the case of Sobolev regularization, $\varphi_{Sobolev}(W)=\mu||\mathcal{D}(W)||^{2}_{F}$ is differentiable, so a term is added to the gradient:
$\varphi_{Sobolev}^{\prime}(W)=2\mu\mathcal{D}^{T}\mathcal{D}(W).$ (20)
Like the other operators, $\mathcal{D}$ can be interpreted as a product with the 2D matrix $D$, since the operator $\mathcal{D}$ is in fact a convolution with a kernel of $[1,-1]$ on the lines, followed by a convolution with a kernel of $[1,-1]^{T}$ on the columns). $\mathcal{D}^{T}$ applies the product with $D^{T}$.
In the code, the operators $\mathcal{R}$ and $\mathcal{D}$ are written as matrix products. The operators $\mathcal{B}_{Y}$ and $\mathcal{S}_{X}$, however, can be dealt with more efficiently by taking advantage of the diagonalization of the convolution operation in the Fourier domain. To do so, we pad the kernels and $Z$ with zeros so they’re the same size (dimension of $Z$ + dimension of the kernel in question). Then, we take their Fourier transform, and convolution becomes a Hadamar product in the Fourier space. The Fourier transform of $Z$ will need to be calculated at each iteration, but the transform of the kernels need only be calculated once.
Putting all this together, we obtain Algorithm 1, which we have dubbed HIFReD: Hyperspectral Image Fusion with Regularized Deconvolution. For simplicity, rebinning is not explicit in the pseudo-code.
1 Input: data ($X$, $Y$), instrument response kernels ($S_{X}$, $B_{Y}$), regularization function ($\varphi$), Low Rank (True/False), rank $r$, regularization parameter $\mu$, number of wavelet scales $S_{2D},S_{1D}$, maximum number of iterations $t_{max}$.
2PWS: Proximal Wavelet Step, defined in Alg. 2.
3Output: $Z$
4InitializationCalculate $\mathcal{F}^{1D}(S_{X}),\mathcal{F}^{2D}(B_{Y})$. $Z\leftarrow X$if Low Rank then
5       $\hat{V}\leftarrow$ eigenvectors of $(Y-\bar{Y})(Y-\bar{Y})^{T}$
6      $V\leftarrow\hat{V}[0:r,.]$
7else
8       $V\leftarrow 1$
[…]
10while $t<t_{max}$ do
11       $W\leftarrow V^{T}Z$$\tilde{W_{t}}=\mathcal{F}^{1D}(\mathcal{F}^{2D}(W_{t}))$. —Gradient Step—if $\varphi(W)=||\mathcal{D}(W)||_{F}^{2}$ then
12             $\tilde{W_{t+1}}\leftarrow\tilde{W_{t}}-\alpha\Big{[}\nabla_{\tilde{W}}\mathcal{L}_{P}(\tilde{W_{t}})-\mu\varphi_{Sobolev}^{\prime}(W_{t})\Big{]}$   Eq. 17, 20.
13      else
14             $\tilde{W_{t+1}}\leftarrow\tilde{W_{t}}-\alpha\Big{[}\nabla_{\tilde{W}}\mathcal{L}_{P}(\tilde{W_{t}})\Big{]}\quad\quad$Eq. 17.
15      $W_{t+1}\leftarrow\mathcal{IF}^{1D}(\mathcal{IF}^{2D}(\tilde{W_{t+1}}))$ —Proximal Step—if $\varphi(W)=||\mathcal{W}_{S}^{2D}(W)||_{1}$ then
16             $W_{t+1}\leftarrow\texttt{PWS}\Big{[}W_{t+1},\mathcal{W}^{2D}_{\mu,S2D}\Big{]}$
17      if $\varphi(W)=||\mathcal{W}_{S}^{2D-1D}(W)||_{1}$ then
18             $W_{t+1}\leftarrow\texttt{PWS}\Big{[}W_{t+1},\mathcal{W}^{2D-1D}_{\mu,(S1D,S2D)}\Big{]}$
19      $Z_{t+1}\leftarrow W_{t+1}V$
20      $Z_{t+1}\leftarrow Z_{t+1}\odot\delta(Z_{t+1}\geq 0)$
21      $t\leftarrow t+1$
return $Z_{t}$
Algorithm 1 HIFReD: Hyperspectral Image Fusion with Regularized Deconvolution
1 input cube $Z$, wavelet transform $\mathcal{W}$, regularization parameter $\mu$, number of wavelet scales $S$$\{w_{c},w_{1},...,w_{S}\}\leftarrow\mathcal{W}(Z)$ $\rho\leftarrow\mu\,\textbf{mad}(w_{1,1})$  (median absolute deviation) for  $i$ in $\{1,...,S\}$ do
2       $w_{i}\leftarrow\textbf{prox}_{l1}(w_{i},\rho)$  Eq. 16
$\hat{Z}\leftarrow w_{c,c}+\sum_{i=1}^{S}w_{i}$ Return( $\hat{Z}$ )
Algorithm 2 PWS: Proximal Wavelet Step
B Hyperparameters
For the choice of hyperparameters (namely, the regularizing factor $\mu$, and the gradient step $\alpha$, and the rank $r$), we need to take a heuristic approach. Indeed, the Poisson likelihood is not globally Lipschitz-continuous, so we cannot analytically calculate an optimal gradient step, and there is no a priori best choice for the regularization parameters either. Further, we found that one set of hyperparameters that worked best for one toy model was not necessarily ideal for another model. The photon count, level of spectral variability, shape of the effective area, and more, are all factors that will impact the choice of best parameters. It would be very costly to compute every single combination of hyperparameters to find the best suited for every single model, but below, we describe some guidelines to help pick suitable parameters, if not the exact optimal value.
The gradient step should be taken as high as possible, for a faster rate of convergence, but if it is taken too high, the likelihood with diverge dramatically. This is easily diagnosed, seen in huge jumps in the cost function, when it should decrease monotonously. This usually happens in early iterations, so we checked the value of the likelihood periodically at the start (first 1000 iterations, taking less than an hour), then stopped if it became pathological. If it was, we then lowered the gradient step and tried again with that parameter.
For the $\mu$ parameter of the Sobolev regularization, for the Gaussian toy model, we tested values between 0.01 and 0.1, and found the best value to be 0.03. This value turned out well suited for the other toy models as well. A too high value of this parameter is detectable by an oversmoothing of spatial features, while a too low value will be recognized in the presence of visible noise on spatial slices.
For the rank $r$, we tested values between 5 and 50, and selected 10. The rank has a similar impact as $\mu$, but on the spectral features. A lower rank might oversmooth and lead to a biased result, while a higher rank might not be enough of a denoiser. On top of that, the rank will impact the algorithm’s computational complexity, with a lower rank implying less dimensions and thus a faster algorihtm.
[…]
$\mu=\frac{(3-0.5)}{1+e^{\tilde{Z}}}+0.5,$ (21)
which will make $\mu(i,j,k)$ close to 1.25 when $|C_{Z}|$ is close to $\bar{|C_{Z}|}$, and go towards 0.5 for small values of $|C_{Z}|$, and towards 3 for high values of $|C_{Z}|$. This simple scheme is however ineffective in cases when the toy-model is not very sparse in the wavelet domain, as was the case for the realistic model. In that case, it was better to pick a low, constant value of $\mu$, such as 0.5.
The last hyperparameter to consider is $t_{max}$, the maximum number of iterations. The user should check the cost function curve at the end, to see whether optimization has converged. If not, they may begin another batch of iterations, using the last result as first guess. An automatic stopping criterion could easily be implemented, with iterations stopping at iteration $t$ once $\Big{|}\frac{\mathcal{L}_{P}(Z_{t})-\mathcal{L}_{P}(Z_{t+1})}{\mathcal{L}_{P}(Z_{t})}\Big{|}>\epsilon$ for a small $\epsilon$, but the results in this article were obtained with a manually set $t_{max}$.
3 Results on simulated data
In this section, we present the results of our method on toy models derived from hydrodynamics simulations of a supernova remnant. Section 3.1 presents our three considered toy models: the Gaussian model, the Gaussian model with rebinning, and the realistic model with rebinning. Section 3.2 then presents our results for each, using the three regularization schemes described in section 2.3.
Figure 3: Comparison of the four PSFs used to generate the toy models. On the left, the Gaussian approximation of a XRISM-like and XMM-Newton-like instrument. On the right, the actual PSFs of XRISM/Resolve and XMM-Newton/EPIC. All PSFs are shown on images of size 4’50”$\times$4’50”. Colors are displayed in square-root scale.
Figure 4: Comparison of the four spectral responses used to generate the toy models. The responses for the XMM-Newton-like instrument is on the right. The ones for the XRISM-like is on the left. The full orange line is for the Gaussian approximations, and the dashed blue line is for the realistic instrument responses.
3.1 Toy Models
Figure 5: Average angular distance between neighbouring pixels on a window of neighbouring pixels, given by Equation 22. Higher values indicate a location of high spectral variability between pixels.
The toy models we constructed were intended to resemble the case study of supernova remnants, such as Cassiopeia A. To do so, we used the 3D hydrodynamics simulations of Orlando et al. (2016). For each voxel of the 3D simulation, a synthetic X-ray spectra was generated using the non-equilibrium ionization Xspec model NEI [^9] https://heasarc.gsfc.nasa.gov/docs/xanadu/xspec/manual/node200.html. In the numerical simulation dataset, the temperature kT ranges from 0.2 keV to 3 keV, the ionization timescale $log(\tau)$ from 8 to 11.5 cm-3 s and the velocity along the line of sight $V_{\rm z}$ from -7000 to 7000 km s-1. The absorption along the line of sight was fixed to $n_{H}=0.5\times 10^{22}$cm-2. Then the hypercube (X, Y, Z, E) is projected along the Z axis to emulate a synthetic X-ray observation of a SNR. Due to the projection effect, the spectrum for each pixel can be quite complex as for example double-peaked spectral features associated to the red and blue-shifted half spheres of the SNR or different temperatures and ionization timescales.
Model Gaussian Gaussian (rebinning) Realistic (rebinning)
Data set X Y X Y X Y
Number of pixels 45$\times$45 45$\times$45 15$\times$15 45$\times$45 15$\times$15
Number of spectral channels a) 2500 - b) 2000 60 2400 60 2400
Energy range a) 0.5-1.4 keV - b) 6.2-6.9 keV 0.5-1.4 keV 0.5-1.4 keV
PSF std 10” 30” 10” 30” 11” 33”
Spectral response std 21 eV 2.5 eV 21 eV 2.5 eV 51 eV 1.4 eV
Photon count a) $2.4\times 10^{7}$ b) $3.2\times 10^{7}$ a) $2.3\times 10^{7}$ b) $3.1\times 10^{7}$ $2.7\times 10^{6}$ $2.5\times 10^{6}$ $4.5\times 10^{7}$ $3.0\times 10^{7}$
Table 1: Table summarizing the characteristics of the simulated data sets. The spectral and spatial responses are given in standard deviation (std).
Then, we wished to consider different levels of complexity for our algorithm benchmarking. In particular, we considered three characteristics which can make the problem more complex: the intensity of spectral variability, the inclusion of a rebinning operator (X and Y have different voxel size), and the ratio between the spectral and spatial responses of the two data sets.
The first toy model we considered is dubbed the Gaussian model. The PSFs and spectral responses of each instrument are assumed to be Gaussian, whose parameters are summarized in Table 1, and shown in Fig. 3 and 4. Further, we assumed a flat effective area, and $X$ having the same size as $Y$. With these parameters, we generated one set of $X$ and $Y$ between 0.5 and 1.4 keV, and one between 6.2 and 6.9 keV (around the iron emission lines). The dataset around the iron lines displays more spectral variability, as is demonstrated more formally in Fig. 5. It shows the average angular distance $\theta$ between neighbouring pixels for each models. For a given pixel’s spectra $Z_{k}$, it is calculated as such:
$\theta\big{(}Z_{k}\big{)}=\sum_{Z_{k^{\prime}}\in\mathcal{N}}\arccos\Big{(}\frac{<Z_{k},Z_{k^{\prime}}>}{\sqrt{(}||Z_{k}||_{2}^{2}*||Z_{k^{\prime}}||_{2}^{2})}\Big{)}/\text{card}(\mathcal{N}),$ (22)
where $\mathcal{N}$ is the set of $Z_{k}$’s neighbourhood, i.e. the 8 closest pixels to $Z_{k}$. This angular distance gives information about the spectral variations across the remnant. As we see, the Gaussian model around the iron line has the strongest variations.
Then, we considered the same Gaussian model (same PSF/spectral response, same flat effective area) between 0.5 and 1.4 keV, but this time with $X$ and $Y$ having different voxel sizes. $X$ is of shape ($45\times 45\times 60$), with square pixels 10” wide and energy channels 14 eV wide, and $Y$ of shape ($15\times 15\times 2400$), with square pixels 30” wide and energy channels 0.36 eV wide. This allows us to see the impact of the rebinning operator on our algorithm.
Finally, we wished to have a more realistic model, which would mimic the case when we try to fuse data from XMM-Newton/EPIC and XRISM/Resolve. We chose PSFs and spectral responses meant to be as close to possible to that of the instruments in question (see Fig. 3 and 4), and likewise for the effective areas. The energy range was (0.5-1.4 keV), and the same dimensions and voxel size as the Gaussian model with rebinning.
On all these case studies, we generated toy-models with a photon count around $10^{7}$. We then applied Poisson noise. Some sample spectra and 2D slices are shown in Fig. 6.
3.2 Results
This section presents our results obtained on the four toy models. The following abbreviations will be used for the three regularization methods: W2D1D for the Wavelet 2D-1D regularization, LRS for low rank approximation with Sobolev regularization, and LRW2D for low rank approximation with Wavelet 2D regularization. We present our results in the form of sample pixel fits, amplitude maps, error histograms, error maps, and scalar metrics.
First, the reconstructed spectra obtained on two sample pixels are shown in Figure 7. Once again, for visual clarity, we only show the method with wavelet 2D-1D sparsity, and the low rank with Sobolev regularization. These results allow us to draw several conclusions. First of all, the Gaussian model around the iron line (c, d) is difficult to reconstruct for both methods, but the LRS method appears more biased. Second, we see that the W2D1D method results in smoother spectra, but this can come lead to more bias, as seen in the Gaussian model with rebinning (e,f). This occurs when the underlying assumption of sparsity in the wavelet domain is not appropriate.
The amplitude maps are displayed in Fig. 8. Overall, for all regularizations, we see that the method is able to reconstruct the general spatial features of the remnant. For the Gaussian model with rebinning, we do notice some amount of leakage on the edges of the remnant for all regularizations, particularly on the western side. For the Gaussian model around 6 keV (b), we also see that the Sobolev regularization (LRS) has oversmoothed some of the features — this effect is visible in the other toy models but more subtle. Finally, in models with rebinning (c, d), the low rank with 2D wavelet sparsity seems to smooth out finer features and give a biased result.
In Fig. 9, we present the map of the spectral angle mapper, an error metric defined as (for a spectra $Z_{i}$) :
$SAM(Z_{i})=\arccos\Bigg{(}\frac{\langle Z_{i},\hat{Z}_{i}\rangle}{||Z_{i}||_{2}||\hat{Z}_{i}||_{2}}\Bigg{)}.$ (23)
The SAM is best if low (close to 0), and aims to capture spectral distortions. In Fig. 9, we observe some leakage on the low-count pixels at the edges of the remnant, present for all methods, but more so for the W2D1D. (a) suggests that the wavelet regularization presented slightly more spatial distorsion for the simplest model (Gaussian model with low spectral variability), while it succeeded better in (b), for the Gaussian model with high spectral variability. In (b), the higher SAM pixels in LRS and LRW2D correspond to some local features, in areas of high spectral variability (see the middle of Fig. 5). Notice, for instance, the cluster of bright spots above the center of the remnant. In (c), the low rank with Sobolev appears to be doing better in the centre. All models display a high error close to the bright spot south of the remnant. For (d), the realistic model, LRS also fares better in the centre of the remnant. While all three methods seem to struggle in the bright area in the South of a remnant, LRW2D and W2D1D do the worst, which could suggest the the Sobolev regularization is better at handling outliers. Since this bright spot is also seen as a problem in the (c) model, this area of high flux and high spatial variability seems especially difficult to reconstruct in models with rebinning.
In Fig. 10, we display the relative error maps averaged per pixel. We draw similar conclusions about the Gaussian model without rebinning. The difficulties for low-rank methods to reconstruct high spectral variabilities areas are once again seen in (b). For (c), we notice again the leakage on the edges, with the amplitude being judged too high in red areas, and too low in blue areas. There is a general bias on all models that seems to overestimate the North-West part of the remnant, while underestimating the South-East part. This is likely due to the rebinning, since it is also present in (d). Further, in (d), we notice that W2D1D and LRW2D underestimated the amplitude of the bright spot in the South of the remnant, while LRS was not as biased. Overall, W2D1D is more biased on the southern bright outlier than the other methods.
Finally, we measured scalar error metrics. All results can be found in Table 2. The best results are highlighted in bold. As Zhang (2008) showed for the case of fusion between panchromatic and MS images, scalar metrics do not necessarily agree when comparing the efficacy of fusion methods, and are not always easily interpretable. Thus, we present four metrics, to provide a more complete comparison. The definition of the metrics’ definition can be found in the Appendix, section A. Namely, we calculate the Normalised Mean Squared Error (NMSE), the average Spectral Angle Mapper (aSAM), average complementary Structural Similarity (acSSIM) index, and the Relative Dimensionless Global Error (ERGAS).
(a) Gaussian model between 0.5-1.4 keV.
(b) Gaussian model between 6.2-6.9 keV.
(c) Gaussian model with rebinning between 0.5-1.4 keV.
(d) Realistic model between 0.5-1.4 keV. [FA: this inculdes rebinning right. Add ”including rebinning” if so.]
Figure 6: Example of spectra and 2D slices from the four toy model data sets. In black, the ground truth $\hat{Z}$. In purple and orange, the two data sets, $X$ with the higher spatial resolution, and $Y$ with the higher spectral resolution.
(a) Gaussian model between 0.5-1.4 keV, pixel (22,22).
(b) Gaussian model between 0.5-1.4 keV, pixel (11,31).
(c) Gaussian model between 6.2-6.9 keV, pixel (22,22).
(d) Gaussian model between 6.2-6.9 keV, pixel (11,31).
(e) Gaussian model with rebinning, pixel (22,22).
(f) Gaussian model with rebinning, pixel (11,31).
(g) Realistic model, pixel (22,22).
(h) Realistic model, pixel (11,31).
Figure 7: Example of how the fusion output compares to the ground truth for two pixels. On the left, pixel (22,22), which is lower in brightness than pixel (11,31), on the right.
Looking at the obtained values in Table 2, we see that the four metrics rarely draw the same conclusion, except in the case of the Gaussian model around 6 keV, i.e. the model with the highest spectral variability between pixels. This allows us to draw a similar conclusion as the error maps: the methods using a low rank approximation have more trouble accounting for high spectral variability. This may be due to the fact that the low rank approximation is learnt by performing PCA on $Y$, the data set which does have the best spectral resolution, but its poor spatial resolution will smooth out spectral variability between pixels. If the low rank could be learnt on $Z$, this would not be a problem, but of course that isn’t possible.
The results on the Gaussian model around 1 keV have all three methods be the best by at least one metric, so it is hard to draw conclusion. In this case, low rankedness does seem to lower spectral distorsion (looking at the aSAM and the ERGAS). For the Gaussian model around 1 keV with rebinning, the low rank methods were more successful, but the difference is slim, and the bright southern spot we noticed on the maps likely plays a part in the the W2D1D being a little less successful. We also observe this for the realistic model with rebinning. Overall, it appears that the low rank methods were best at dealing with outliers, especially the low rank with Sobolev, but they were not as appropriate for models with high spectral variability.
Convergence took between 30k and 100k iterations, depending on the toy model. On average, one iteration of the algorithm took around 3s with the method with 2D-1D wavelet regularization, resulting in computation time between 25h-83h. With the low rank approximation with Sobolev, each iteration took around 2s, resulting in computation time between 15h-55h. Low rank with Wavelet 2D was around the same as the latter. The reduction in dimension of the low-rankedness allows these methods to be much faster.
Gaussian (0.5-1.4 keV)
Regularization NMSE aSAM acSSIM ERGAS
W2D1D 21.36 0.571 0.017 0.220
LRS 20.97 0.557 0.011 0.116
LRW2D 21.01 0.560 0.012 0.114
Gaussian (6.2-6.9 keV)
NMSE aSAM acSSIM ERGAS
W2D1D 18.67 0.599 0.005 0.229
LRS 14.16 0.600 0.011 0.343
LRW2D 16.57 0.609 0.010 0.248
Gaussian with Rebinning (0.5-1.4 keV)
NMSE aSAM acSSIM ERGAS
W2D1D 10.36 0.691 0.048 0.422
LRS 11.01 0.670 0.045 0.341
LRW2D 10.34 0.690 0.054 0.317
Realistic with Rebinning (0.5-1.4 keV)
NMSE aSAM acSSIM ERGAS
W2D1D 9.956 0.700 0.032 0.608
LRS 10.16 0.691 0.031 0.358
LRW2D 9.523 0.712 0.038 0.351
Table 2: Table of error metrics, comparing the results obtained by the three potential regularizors for the four toy models. The NMSE (Equation 24) is better if high, while the aSAM (Equation 25) and acSSIM (Equation 26) are better if low. The best values are marked in bold.
(a) Gaussian model between 0.5-1.4 keV.
(b) Gaussian model between 6.2-6.9 keV.
(c) Gaussian model with rebinning between 0.5-1.4 keV.
(d) Realistic model between 0.5-1.4 keV.
Figure 8: Amplitude maps (i.e. sum of counts per pixel over every spectral channel) obtained by the three regularizers and compared to the ground truth.
(a) Gaussian model between 0.5-1.4 keV.
(b) Gaussian model between 6.2-6.9 keV.
(c) Gaussian model with rebinning between 0.5-1.4 keV.
(d) Realistic model between 0.5-1.4 keV.
Figure 9: Maps of the SAM (Equation 23) for the four toy models, and the three regularizations. A higher SAM means a worse spectral reconstruction.
(a) Gaussian model between 0.5-1.4 keV.
(b) Gaussian model between 6.2-6.9 keV.
(c) Gaussian model with rebinning between 0.5-1.4 keV.
(d) Realistic model between 0.5-1.4 keV.
Figure 10: Maps of the relative error: $Err(Z)=100(Z-\hat{Z})/\hat{Z}$, for the four toy models, and the three regularizations. A higher absolute error means a worse spectral reconstruction.
4 Discussion
Overall, the Wavelet 2D-1D regularization was more successful than low rank approximations to deal with strong spectral variability across the image. However, the assumption of wavelet sparsity is not necessarily ideal to account for certain structures, such as in the presence of singular bright spots like the one in the southern part of the remnant, and this was especially difficult to treat with the inclusion of a rebinning operator. Nonetheless, we found that for all three regularization methods, the fusion results were satisfactory given the difficulty of the problem, but they still contained some bias.
The low rank approximation, while ill-suited for supernova remnants that display strong spectral variability due to the variations of their physical parameters, was found to accelerate the algorithm by half, thanks to the dimension reduction. Since the data sets are large and the algorithm takes long to converge, this is a useful attribute.
Thus, it would be ideal to find a spectral regularization less biased than the wavelet sparsity or the low rank approximation, but which nonetheless includes a reduction in dimensions. This may be done by exploring a low rank approximation more refined than one learnt using PCA on the high spectral resolution data set. We could also explore acceleration methods, such as algorithm unrolling Fahes et al. (2022).
Alternatively, a promising avenue would be to include physical information. Indeed, while having a blind fusion result is useful (as it makes no assumptions on the model), in order to obtain physical parameter maps, a model will need to be fitted to the result. In the case of supernova remnants, it would in fact be more than one model, since the emission comes from a mixture composed of thermal radiations from the shocked interstellar media, and synchrotron radiation due to particle acceleration.
An algorithm for source separation that includes spectral variations was developed in Lascar, J. et al. (2024). Our future work will focus on combining this approach to HIFReD, and thus develop a method capable of simultaneously fuse data sets from two generations of telescopes and perform unmixing with a physical spectral model.
5 Conclusions
We have developed an algorithm for fusing hyperspectral images taken by two generations of X-ray telescopes with complementary spatial and spectral resolutions. Our method simultaneously deconvolves the instruments’ spatial and spectral responses, in order to return to the best resolutions of the two telescopes.
We implemented three regularisation schemes, the first being a term to promote sparsity under a 2D-1D wavelet transform, the second being a low rank approximation with Sobolev regularization (inspired by Guilloteau et al., 2020), and the third being a low rank approximation with a 2D wavelet sparsity term.
We tested our method on five toy-models, constructed from hydrodynamic simulations of supernova remnants by Orlando et al. (2016) convolved with an increasing complexity of instrumental response. Spatial features were well reconstructed, apart from a slight leakage in the bright shell in the toy-model with rebinning. The reconstruction of spectral features depended on the regularizations and model.
We found that while low rank methods accelerated the algorithm by half, they struggled to account for the strong spectral variability present around the iron line. The 2D-1D wavelet method accounted better for this variability, but it was somewhat biased in reconstructing the realistic model.
In conclusion, we believe that hyperspectral fusion shows promise and would greatly benefit the analysis of XRISM/Resolve data, though it would be useful to explore other regularization terms more suited to the variabilities of the data. In particular, we intend to combine fusion with a spectral physical model and source separation algorithm in future works. While the purpose of this article was a proof of concept, we are now actively working on applying this method to real data.
Acknowledgements.
The research leading to these results has received funding from the European Union’s Horizon 2020 Programme under the AHEAD2020 project (grant agreement n. 871158). This work was supported by CNES, focused on methodology for X-ray analysis. This work was supported by the Action Thématique Hautes Energies of CNRS/INSU with INP and IN2P3, co-funded by CEA and CNES. We thank S. Orlando for kindly providing the simulation from Orlando et al. (2016) to generate our toy model.
Appendix A Metrics definition
In this section, we define the scalar metrics calculated in Table 2. First, the Normalised Mean Squared Error (NMSE) is defined as such (where $Z_{i}$ is the value of $Z$ at the voxel $i$):
$NMSE(\hat{Z},Z)=10\log_{10}\frac{\sum_{i=1}^{LMN}(\hat{Z}_{i}^{2})}{\sum_{i=1}^{LMN}\Big{(}(\hat{Z}_{i}-Z_{i})^{2}\Big{)}},$ (24)
By this definition, a higher NMSE implies a result of better quality. Then, the average Spectral Angle Mapper (aSAM) is defined as:
$aSAM(\hat{Z},Z)=\frac{1}{L}\sum_{i}^{MN}\arccos\Bigg{(}\frac{\langle Z_{i},\hat{Z}_{i}\rangle}{||Z_{i}||_{2}||\hat{Z}_{i}||_{2}}\Bigg{)},$ (25)
where $Z_{i}$ is the spectra for the pixel $i$. A lower aSAM implies a result of better quality, and an ideal value is 0.
The average complementary Structural Similarity (SSIM) index, is:
$acSSIM(\hat{Z},Z)=1-\frac{1}{L}\sum_{k=1}^{L}\text{SSIM}(\hat{Z}_{k},Z_{k}),$ (26)
where SSIM is the Structural Similarity index, as defined by Wang et al. (2004). The aim of the acSSIM is to measure the capacity to reconstruct local image structure, luminance, and contrast. The lower the acSSIM, the better the result. An ideal value is 0.
Finally, the Relative Dimensionless Global Error (ERGAS), is defined as:
$\text{ERGAS}(\hat{Z},Z,d)=100d\sqrt{\frac{1}{L}\sum_{k=1}^{L}\Bigg{(}\frac{||\hat{Z}_{k}-Z_{k}||_{F}}{\mu_{k}\sqrt{MN}}\Bigg{)}^{2}},$ (27)
where $d$ is the ratio between the spatial resolution of $X$ and the spatial resolution of $Y$, and $\mu_{k}$ is the mean of $\hat{Z}_{k}$. The lower the ERGAS, the better, and an ideal value is 0.